# Probing Dark Matter Substructures with Free-Form Modelling: A Case Study of the ‘Jackpot’ Strong Lens

### v1

@@ Abstract
Characterising the population and internal structure of sub-galactic halos is critical for constraining the nature of dark matter. These halos can be detected near galaxies that act as strong gravitational lenses with extended arcs, as they perturb the shapes of the arcs. However, this method is subject to false-positive detections and systematic uncertainties: particularly degeneracies between an individual halo and larger-scale asymmetries in the distribution of lens mass. We present a new free-form lens modelling code, developed within the framework of the open-source software PyAutoLens, to address these challenges. Our method models mass perturbations that cannot be captured by parametric models as pixelized potential corrections and suppresses unphysical solutions via a Matérn regularisation scheme that is inspired by Gaussian process regression. This approach enables the recovery of diverse mass perturbations, including subhalos, line-of-sight halos, external shear, and multipole components that represent the complex angular mass distribution of the lens galaxy, such as boxiness/diskiness. Additionally, our fully Bayesian framework objectively infers hyperparameters associated with the regularisation of pixelized sources and potential corrections, eliminating the need for manual fine-tuning. By applying our code to the well-known ‘Jackpot’ lens system, SLACS0946+1006, we robustly detect a highly concentrated subhalo that challenges the standard cold dark matter model. This study represents the first attempt to independently reveal the mass distribution of a subhalo using a fully free-form approach.
@@
pubyear: 2025pagerange: Probing Dark Matter Substructures with Free-Form Modelling: A Case Study of the ‘Jackpot’ Strong LensB
@@ 1 Introduction
The standard cosmological model has achieved remarkable success (Planck Collaboration et al., 2020), as it explains numerous observations, such as the rotation curves of disk galaxies (Rubin, 1991), the clustering properties of galaxies (Springel et al., 2005), the properties of the hot gas in galaxy clusters (Ettori et al., 2013), and the cosmic microwave background (Planck Collaboration et al., 2014). The standard cosmological model assumes that approximately 85 percent of the matter in the universe is the so-called "dark matter" (Planck Collaboration et al., 2020). Dark matter neither emits light nor interacts electromagnetically with baryonic matter. Instead, its presence is inferred solely from gravitational effects (Arbey & Mahmoudi, 2021). The fundamental nature of dark matter remains elusive; it may be primordial black holes (Carr et al., 2016) or some undiscovered particles beyond the standard model of particle physics (Feng, 2010). Since different types of dark matter produce distinct cosmological structures, such as dark matter halos and subhalos, astronomical observations can serve as an effective means to constrain the properties of dark matter (Frenk & White, 2012). The popular probes currently used to infer the properties of dark matter include stellar streams in the local group (Ibata et al., 2001; Bonaca et al., 2019), the Lyman-alpha forest (Seljak et al., 2006), the luminosity function of Milky Way satellite galaxies (Nadler et al., 2021; Newton et al., 2021), and strong gravitational lensing (Vegetti et al., 2023). Among these, the strong gravitational lensing technique offers a distinct advantage, as it can infer the properties of dark matter solely based on the gravitational distortion effects of dark matter structures, without relying on complex baryonic physics, and it can be applied to higher redshifts.
@@
Strong gravitational lensing occurs when light from a background source is significantly distorted by a massive foreground deflector, such as a galaxy or a galaxy cluster, forming multiple images or extended arcs (Schneider et al., 1992; Meylan et al., 2006; Meneghetti, 2022). Smaller dark matter halos within the main deflector (subhalos) or along the line of sight (“LOS-halos”, Li et al., 2017; Despali et al., 2018; He et al., 2022b) can further perturb the lensing signal, creating observable anomalies in lensed images, time delays, and flux ratios between multiple images. These anomalies serve as valuable probes for constraining the properties of dark matter (Vegetti & Koopmans, 2009b; Li et al., 2016; He et al., 2022a; Keeton & Moustakas, 2009; Liao et al., 2018; Mao & Schneider, 1998; Birrer et al., 2017; Gilman et al., 2018; Hsueh et al., 2020). Current time-delay measurements, typically with uncertainties on the order of a few days (Millon et al., 2020), lack the precision required to detect subhalo-induced signals, which are often on the order of fractions of a day. Flux ratio anomalies, theoretically most sensitive to subhalo perturbations due to their direct dependence on the second derivative of the lensing potential, generally cannot precisely constrain the properties (e.g., mass and position) of individual subhalos (e.g. Nierenberg et al., 2014). This limitation arises from the sparsity of observational data (usually restricted to positions and fluxes of lensed images), which does not provide sufficient constraints to resolve degeneracies either among subhalo internal properties (e.g., mass-position degeneracy) or between the subhalo and the main lens galaxy (Fadely & Keeton, 2012; Hsueh et al., 2020; Cohen et al., 2024). Consequently, subhalo properties inferred from flux ratio anomalies are typically derived statistically—finding multiple feasible solutions across a range of subhalo masses and positions—instead of precise individual measurements (Dalal & Kochanek, 2002; Hsueh et al., 2020; Gilman et al., 2020). In contrast, the brightness distribution along distorted lensed arcs, typically comprising several thousand pixels, can effectively break the degeneracies encountered by flux ratio anomalies. Thus, analyzing extended arcs offers a more reliable method for precisely measuring the properties of individual subhalos through strong gravitational lensing (e.g., Vegetti et al., 2010, 2012; Hezaveh et al., 2016; Ritondale et al., 2019; Nightingale et al., 2024).
@@
Subhalo detections using distorted lensed arcs have been widely employed to constrain dark matter properties (e.g., Vegetti et al., 2014, 2018; Enzi et al., 2021). Since the distortions induced by small-mass subhalos are subtle, sophisticated lens modelling methods are essential for extracting these weak signals. Currently, two primary approaches are employed: parametric and free-form methods. The parametric method uses explicit parametric models to describe the mass distribution of both the main lens galaxy and the presumed subhalo. In this approach, a subhalo is considered detected if a model including a subhalo provides a significantly better fit to the observational data, as measured by an increase in Bayesian evidence, compared to a model without a subhalo (Vegetti & Koopmans, 2009a; Ritondale et al., 2019; Nightingale et al., 2024). In contrast, the free-form method approximates the mass distribution of the main lens galaxy using a smooth parametric model while treating perturbative signals not captured by the parametric model as pixelized linear corrections to the lensing potential [^1] This free-form approach to the pixelized perturbative lensing potential is widely known as the potential correction method. (Koopmans, 2005; Suyu & Blandford, 2006; Suyu et al., 2009; Vegetti & Koopmans, 2009a; Ritondale et al., 2019; Vernardos & Koopmans, 2022; Galan et al., 2022; Rüstig et al., 2024). Subhalos appear as localised positive mass clumps in the resulting mass perturbation map. While the parametric method is effective at quantifying the properties of subhalo perturbers, it is prone to false positives, where a subhalo is incorrectly inferred to exist (Ritondale et al., 2019). This issue stems from systematic errors in the parametric model, particularly degeneracies between the mass distribution of the main lens galaxy and the subhalo. If the main lens galaxy’s mass model is oversimplified, a false-positive subhalo signal may arise to compensate for missing complexity in the lens galaxy (He et al., 2023; Nightingale et al., 2024; O’Riordan & Vegetti, 2024). Although adding complexity to the lens mass model can test the robustness of subhalo detections (Lange et al., 2024), it is often unclear whether the added complexity is sufficient to eliminate all false positives. Consequently, subhalo detection claims based solely on parametric methods may be met with scepticism. To address this concern, Vegetti et al. (2014) advocated for free-form approaches to validate subhalo detections obtained using parametric methods.
@@
Regularisation, or smoothing, is essential for obtaining physically meaningful solutions in free-form approaches (Wallington et al., 1996; Warren & Dye, 2003). The classical potential correction method begins with an initial estimate of the main lens mass distribution and the source brightness distribution. A strong regularisation strength is empirically applied to the pixelized lensing potential perturbations, producing a minor correction to the lensing potential (Koopmans, 2005; Vegetti & Koopmans, 2009a). This correction is subsequently used to refine both the lens mass and source light models, initiating a new iteration of potential correction. The process is repeated iteratively until convergence is achieved, defined as the point where additional corrections no longer result in significant improvements to the model fit (Koopmans, 2005). A major limitation of this iterative algorithm is its vulnerability to instability: artefacts caused by an inappropriate choice of regularisation strength can accumulate over iterations, ultimately compromising the modelling process. Vernardos & Koopmans (2022) introduced a framework to objectively determine the regularisation strength for pixelized lensing potential perturbations by maximising Bayesian evidence and directly solving for the potential correction, thereby avoiding the need for unstable iterative algorithms. However, the Gaussian and Exponential regularisation techniques used in their study are more effective for extended perturbations but perform poorly when applied to localised features, such as those caused by subhalos [^2] Further demonstration of this limitation is provided in Section 4.1..
@@
In this work, we present a new free-form lens modelling code built upon the open-source software PyAutoLens (Nightingale & Dye, 2015; Nightingale et al., 2021, 2018). By utilising a generalised Matérn regularisation (Rasmussen & Williams, 2006) and determining the regularisation parameters objectively through the Bayesian evidence, our code can automatically recover diverse forms of mass perturbation signals without requiring manual fine-tuning. To validate our method, we first test it on mock lensing systems, demonstrating its robustness in recovering various forms of lensing potential perturbations commonly encountered in strong lensing studies. We then apply our method to the strong lens system SLACS0946+1006 (Gavazzi et al., 2008; Smith & Collett, 2021), also known as the “Jackpot” lens, and characterise the mass perturbations in this system using a fully free-form approach.
This paper is organised as follows. In Section 2, we describe the methodology for the potential correction. Section 3 describes the datasets employed in this study, including mock datasets and the real “Jackpot” lens system (SLACS0946+1006). In Section 4, we first evaluate the performance of our method in quantitatively recovering mass perturbation signals using mock datasets and then present a free-form analysis of the real lens system SLACS0946+1006. The strengths, limitations, and future prospects of our method are discussed in Section 5, and a concise summary is provided in Section 6. Unless otherwise specified, all subsequent references to ‘best-fit’ parameter values in this work refer to the median values of the posterior distribution. Throughout this work, we assume a flat $\Lambda$CDM cosmology with $\Omega_{\rm m}=0.3$, $\Omega_{\Lambda}=0.7$, and $H_{0}=70\,\mathrm{km}\,\mathrm{s}^{-1}\,\mathrm{Mpc}^{-1}$. All codes and relevant data used in this work are publicly available at the following link: https://github.com/caoxiaoyue/lensing_potential_correction.
@@ 2.2 Inverting Only the Potential Perturbation
The primary goal of the potential correction algorithm is to solve the partial differential Equation 5, using the image residual and source gradient from the macro model. A straightforward method for solving this equation is numerical integration along the characteristic curve, as outlined in Suyu & Blandford (2006). However, this approach is effective only when the perturbative lensing potential is small (less than 1%). A more robust alternative for solving equation 5 is to reformulate it as the following matrix expression:
@@
    $\delta\boldsymbol{d}$ is a column vector where each element $\delta d_{j}$ represents the value of image residuals provided by the macro model for each image pixel, where $j=1,\ldots,N_{d}$, and $N_{d}$ denotes the number of image pixels. Similarly, the vector $\boldsymbol{n}$ represents the noise realisation at each image pixel, with values drawn from the multivariate normal distribution $\mathcal{N}(0,\,\mathrm{C}_{\mathrm{D}})$, where $\mathrm{C}_{\mathrm{D}}$ is the covariance matrix.
@@
    $\boldsymbol{B}$ is a square matrix of dimensions $[N_{d},N_{d}]$ constructed to account for the blurring effect of the PSF.
@@
    $\boldsymbol{\beta_{j}^{d}}$ represents the ray-traced source plane position of the $j$-th native image grid (with position denoted by $\boldsymbol{\theta_{j}^{d}}$), according to the macro model. $S$ denotes the source’s brightness distribution. The $x$ and $y$ components of the source’s brightness gradient are given by $\frac{\partial S\left(\boldsymbol{\beta_{j}^{d}}\right)}{\partial\beta_{x}}$ and $\frac{\partial S\left(\boldsymbol{\beta_{j}^{d}}\right)}{\partial\beta_{y}}$.
@@
For convenience, we abbreviate $-\boldsymbol{B}\boldsymbol{D}_{\mathrm{s}}\left(\boldsymbol{s_{\mathrm{M}}}\right)\boldsymbol{C}_{\mathrm{f}}\boldsymbol{D}_{\psi}$ as $\mathbf{L_{\delta\psi}}$. The latter has dimensions of $N_{d}\times N_{p}$.
It is evident that $\delta\boldsymbol{d}$ and $\boldsymbol{\delta\psi}$ are connected via a linear mapping. The most likely solution for $\boldsymbol{\delta\psi}$ can be derived by maximizing the likelihood
where
$=\frac{1}{2}\chi^{2},$
and
The most likely solution is usually degraded by noise and is ill-posed (Warren & Dye, 2003). An additional prior $P(\delta\boldsymbol{\psi}\mid\boldsymbol{g}_{\delta\psi},\boldsymbol{\xi}_{\delta\psi})$ is needed to impose smoothness on $\delta\boldsymbol{\psi}$, a process commonly called “regularisation”. Here, $\boldsymbol{g}_{\delta\psi}$ represents the type of regularisation, and the associated model parameters are denoted by the vector $\boldsymbol{\xi}_{\delta\psi}$. The prior can be expressed as:
where $Z_{\delta\psi}(\boldsymbol{\xi}_{\delta\psi})=\int\mathrm{d}^{N_{p}}(\delta\boldsymbol{\psi})\exp(-E_{\mathrm{\delta\psi}})$ is the normalization factor of the prior probability density function. The function $E_{\delta\psi}$, also known as the regularisation function, typically takes a quadratic form: $E_{\delta\psi}(\delta\boldsymbol{\psi})=\frac{1}{2}\delta\boldsymbol{\psi}^{\mathrm{T}}\mathbf{R_{\delta\psi}}\delta\boldsymbol{\psi}$, where $\mathbf{R_{\delta\psi}}$ represents the Hessian of $E_{\delta\psi}$. This form allows $\delta\boldsymbol{\psi}$ to be directly solved through matrix inversion (Warren & Dye, 2003). For more detailed information on the specific form of the regularisation function, refer to Section 2.4.
This yields:
$\boldsymbol{\delta\psi}_{\mathrm{mp}}=\left(\mathbf{L^{T}_{\delta\psi}}\mathrm{C}_{\mathrm{D}}^{-1}\mathbf{L_{\delta\psi}}+\mathbf{R_{\delta\psi}}\right)^{-1}\mathbf{L^{T}_{\delta\psi}}\mathrm{C}_{\mathrm{D}}^{-1}\delta\boldsymbol{d}.$ (16)
Equation 16 indicates that given the macro model results $\{\delta\boldsymbol{d},\,\mathbf{L^{T}_{\delta\psi}}\}$ and an arbitrary choice of regularisation $\mathbf{R_{\delta\psi}}(\boldsymbol{\xi}_{\delta\psi},\boldsymbol{g}_{\delta\psi})$, the most probable solution $\boldsymbol{\delta\psi}_{\mathrm{mp}}$ can be obtained via linear matrix inversion. The covariance matrix ($\Sigma_{\delta\psi}$), which describes the uncertainty of the most probable solution ($\boldsymbol{\delta\psi}_{\mathrm{mp}}$), is derived by Suyu et al. (2006) and follows:
Here, the evidence term $P(\delta\boldsymbol{d}\mid\mathbf{L}_{\delta\psi},\boldsymbol{g}_{\delta\psi})$ marginalizes over the hyperparameters of the regularisation and can be used to rank different model choices of $\{\mathbf{L}_{\delta\psi},\boldsymbol{g}_{\delta\psi}\}$. Since a non-informative prior is usually taken for $P(\boldsymbol{\xi}_{\delta\psi})$, finding the most probable solution for $\boldsymbol{\xi}_{\delta\psi}$ is equivalent to maximizing $P(\delta\boldsymbol{d}\mid\mathbf{L_{\delta\psi}},\boldsymbol{\xi}_{\delta\psi},\boldsymbol{g}_{\delta\psi})\equiv\mathcal{E}$, whose formula is derived by Suyu et al. (2006) and reorganized by Dye et al. (2008), as follows:
$\operatorname{log}\mathcal{E}=$ $-\frac{\mathrm{N_{d}}}{2}\log(2\pi)-\frac{1}{2}\log(\det\mathrm{C}_{\mathrm{D}})$ (19)
$+\frac{1}{2}\log(\det\mathbf{R}_{\delta\psi})-\frac{1}{2}\delta\boldsymbol{\psi}^{\mathrm{T}}_{\mathrm{mp}}\mathbf{R}_{\delta\psi}\delta\boldsymbol{\psi}_{\mathrm{mp}}$
$-\frac{1}{2}(\mathbf{L}_{\delta\psi}\delta\boldsymbol{\psi}_{\mathrm{mp}}-\delta\boldsymbol{d})^{T}\mathrm{C}_{\mathrm{D}}^{-1}(\mathbf{L}_{\delta\psi}\delta\boldsymbol{\psi}_{\mathrm{mp}}-\delta\boldsymbol{d})$
$-\frac{1}{2}\log[\det(\mathbf{L^{T}_{\delta\psi}}\mathrm{C}_{\mathrm{D}}^{-1}\mathbf{L}_{\delta\psi}+\mathbf{R}_{\delta\psi})].$

### v2

@@ Abstract
Characterising the population and internal structure of sub-galactic halos is critical for constraining the nature of dark matter. These halos can be detected near galaxies that act as strong gravitational lenses with extended arcs, as they perturb the shapes of the arcs. However, this method is subject to false-positive detections and systematic uncertainties, particularly degeneracies between an individual halo and larger-scale asymmetries in the distribution of lens mass. We present a new free-form lens modelling code, developed within the framework of the open-source software PyAutoLens, to address these challenges. Our method models mass perturbations that cannot be captured by parametric models as pixelized potential corrections and suppresses unphysical solutions via a Matérn regularisation scheme that is inspired by Gaussian process regression. This approach enables the recovery of diverse mass perturbations, including subhalos, line-of-sight halos, external shear, and multipole components that represent the complex angular mass distribution of the lens galaxy, such as boxiness/diskiness. Additionally, our fully Bayesian framework objectively infers hyperparameters associated with the regularisation of pixelized sources and potential corrections, eliminating the need for manual fine-tuning. By applying our code to the well-known ‘Jackpot’ lens system, SLACS0946+1006, we robustly detect a highly concentrated subhalo that challenges the standard cold dark matter model. This study represents the first attempt to independently reveal the mass distribution of a subhalo using a fully free-form approach.
@@
pubyear: 2026pagerange: Probing Dark Matter Substructures with Free-Form Modelling: A Case Study of the ‘Jackpot’ Strong LensC
@@ 1 Introduction
The standard cosmological model has achieved remarkable success (Planck Collaboration et al., 2020), as it explains numerous observations, such as the rotation curves of disk galaxies (Rubin, 1991), the clustering properties of galaxies (Springel et al., 2005), the properties of the hot gas in galaxy clusters (Ettori et al., 2013), and the cosmic microwave background (Planck Collaboration et al., 2014). The standard cosmological model assumes that approximately 85 percent of the matter in the universe is the so-called "dark matter" (Planck Collaboration et al., 2020). Dark matter neither emits light nor interacts electromagnetically with baryonic matter. Instead, its presence is inferred solely from gravitational effects (Arbey and Mahmoudi, 2021). The fundamental nature of dark matter remains elusive; it may be primordial black holes (Carr et al., 2016) or some undiscovered particles beyond the standard model of particle physics (Feng, 2010). Since different types of dark matter produce distinct cosmological structures, such as dark matter halos and subhalos, astronomical observations can serve as an effective means to constrain the properties of dark matter (Frenk and White, 2012). The popular probes currently used to infer the properties of dark matter include stellar streams in the local group (Ibata et al., 2001; Bonaca et al., 2019), the Lyman-alpha forest (Seljak et al., 2006), the luminosity function of Milky Way satellite galaxies (Nadler et al., 2021; Newton et al., 2021), and strong gravitational lensing (Vegetti et al., 2023). Among these, the strong gravitational lensing technique offers a distinct advantage, as it can infer the properties of dark matter solely based on the gravitational distortion effects of dark matter structures, without relying on complex baryonic physics, and it can be applied to higher redshifts.
@@
Strong gravitational lensing occurs when light from a background source is significantly distorted by a massive foreground deflector, such as a galaxy or a galaxy cluster, forming multiple images or extended arcs (Schneider et al., 1992; Meylan et al., 2006; Meneghetti, 2022). Smaller dark matter halos within the main deflector (subhalos) or along the line of sight (“LOS-halos”, Li et al., 2017; Despali et al., 2018; He et al., 2022a) can further perturb the lensing signal, creating observable anomalies in lensed images, time delays, and flux ratios between multiple images. These anomalies serve as valuable probes for constraining the properties of dark matter (Vegetti and Koopmans, 2009b; Li et al., 2016; He et al., 2022b; Keeton and Moustakas, 2009; Liao et al., 2018; Mao and Schneider, 1998; Birrer et al., 2017; Gilman et al., 2018; Hsueh et al., 2020). Current time-delay measurements, typically with uncertainties on the order of a few days (Millon et al., 2020), lack the precision required to detect subhalo-induced signals, which are often on the order of fractions of a day. Flux ratio anomalies, theoretically most sensitive to subhalo perturbations due to their direct dependence on the second derivative of the lensing potential, generally cannot precisely constrain the properties (e.g., mass and position) of individual subhalos (e.g. Nierenberg et al., 2014). This limitation arises from the sparsity of observational data (usually restricted to positions and fluxes of lensed images), which does not provide sufficient constraints to resolve degeneracies either among subhalo internal properties (e.g., mass-position degeneracy) or between the subhalo and the main lens galaxy (Fadely and Keeton, 2012; Hsueh et al., 2020; Cohen et al., 2024). Consequently, subhalo properties inferred from flux ratio anomalies are typically derived statistically—finding multiple feasible solutions across a range of subhalo masses and positions—instead of precise individual measurements (Dalal and Kochanek, 2002; Hsueh et al., 2020; Gilman et al., 2020). In contrast, the brightness distribution along distorted lensed arcs, typically comprising several thousand pixels, can effectively break the degeneracies encountered by flux ratio anomalies. Thus, analysing extended arcs offers a more reliable method for precisely measuring the properties of individual subhalos through strong gravitational lensing (e.g., Vegetti et al., 2010, 2012; Hezaveh et al., 2016; Ritondale et al., 2019; Nightingale et al., 2024).
@@
Subhalo detections using distorted lensed arcs have been widely employed to constrain dark matter properties (e.g., Vegetti et al., 2014, 2018; Enzi et al., 2021). Since the distortions induced by small-mass subhalos are subtle, sophisticated lens modelling methods are essential for extracting these weak signals. Currently, two primary approaches are employed: parametric and free-form methods. The parametric method uses explicit parametric models to describe the mass distribution of both the main lens galaxy and the presumed subhalo. In this approach, a subhalo is considered detected if a model including a subhalo provides a significantly better fit to the observational data, as measured by an increase in Bayesian evidence, compared to a model without a subhalo (Vegetti and Koopmans, 2009a; Ritondale et al., 2019; Nightingale et al., 2024). In contrast, the free-form method approximates the mass distribution of the main lens galaxy using a smooth parametric model while treating perturbative signals not captured by the parametric model as pixelized linear corrections to the lensing potential [^1] This free-form approach to the pixelized perturbative lensing potential is widely known as the potential correction method. Other types of free-form approaches have also been employed in strong lensing analyses, which represent the mass (instead of the lensing potential) on a grid of pixels (e.g. Saha and Williams, 1997) or as a linear combination of basis functions (e.g. Liesenborgs et al., 2006; Diego et al., 2007). (Koopmans, 2005; Suyu and Blandford, 2006; Suyu et al., 2009; Vegetti and Koopmans, 2009a; Ritondale et al., 2019; Vernardos and Koopmans, 2022; Galan et al., 2022; Rüstig et al., 2024). Subhalos appear as localised positive mass clumps in the resulting mass perturbation map. While the parametric method is effective at quantifying the properties of subhalo perturbers, it is prone to false positives, where a subhalo is incorrectly inferred to exist (Ritondale et al., 2019). This issue stems from systematic errors in the parametric model, particularly degeneracies between the mass distribution of the main lens galaxy and the subhalo. If the main lens galaxy’s mass model is oversimplified, a false-positive subhalo signal may arise to compensate for missing complexity in the lens galaxy (He et al., 2023; Nightingale et al., 2024; O’Riordan and Vegetti, 2024). Although adding complexity to the lens mass model can test the robustness of subhalo detections (Lange et al., 2024), it is often unclear whether the added complexity is sufficient to eliminate all false positives. Consequently, subhalo detection claims based solely on parametric methods may be met with scepticism. To address this concern, Vegetti et al. (2014) advocated for free-form approaches to validate subhalo detections obtained using parametric methods.
@@
Regularisation, or smoothing, is essential for obtaining physically meaningful solutions in free-form approaches (Wallington et al., 1996; Warren and Dye, 2003). The classical potential correction method begins with an initial estimate of the main lens mass distribution and the source brightness distribution. A strong regularisation strength is empirically applied to the pixelized lensing potential perturbations, producing a minor correction to the lensing potential (Koopmans, 2005; Vegetti and Koopmans, 2009a). This correction is subsequently used to refine both the lens mass and source light models, initiating a new iteration of potential correction. The process is repeated iteratively until convergence is achieved, defined as the point where additional corrections no longer result in significant improvements to the model fit (Koopmans, 2005). A major limitation of this iterative algorithm is its vulnerability to instability: artefacts caused by an inappropriate choice of regularisation strength can accumulate over iterations, ultimately compromising the modelling process. Vernardos and Koopmans (2022) introduced a framework to objectively determine the regularisation strength for pixelized lensing potential perturbations by maximising Bayesian evidence and directly solving for the potential correction, thereby avoiding the need for unstable iterative algorithms. However, the Gaussian and Exponential regularisation techniques used in their study are more effective for extended perturbations but perform poorly when applied to localised features, such as those caused by subhalos [^2] Further demonstration of this limitation is provided in Section 4.1..
@@
The Matérn kernel is a widely used covariance function that imposes smoothness constraints in free-form reconstruction via Gaussian process regression. This Matérn covariance kernel, or regularisation, has recently been applied to pixelised source reconstruction in strong gravitational lensing analyses (Galan et al., 2024; Enzi et al., 2024; Rüstig et al., 2024). Whilst Vernardos and Koopmans (2022) primarily investigated the advantages of Gaussian and exponential forms of regularisation—special cases of the more general Matérn kernel—for reconstructing pixelised sources and mass models, they also proposed that the Matérn kernel merits further exploration. In this work, we present a new free-form lens modelling code built upon the open-source software PyAutoLens (Nightingale and Dye, 2015; Nightingale et al., 2021, 2018). By employing a generalised Matérn regularisation for the pixelised mass model (in the form of potential corrections) (Rasmussen and Williams, 2006) and determining the regularisation parameters objectively through the Bayesian evidence, our code automatically recovers diverse forms of mass perturbation signals without requiring manual fine-tuning. To validate our method, we first test it on mock lensing systems, demonstrating its robustness in recovering various forms of lensing potential perturbations commonly encountered in strong lensing studies. We then apply our method to the strong lens system SLACS0946+1006 (Gavazzi et al., 2008; Smith and Collett, 2021), also known as the ‘Jackpot’ lens, and characterise the mass perturbations in this system using a fully free-form approach.
[…]
@@ 2.2 Inverting Only the Potential Perturbation
The primary goal of the potential correction algorithm is to solve the partial differential Equation 5, using the image residual and source gradient from the macro model. A straightforward method for solving this equation is numerical integration along the characteristic curve, as outlined in Suyu and Blandford (2006). However, this approach is effective only when the perturbative lensing potential is small (less than 1%). A more robust alternative for solving equation 5 is to reformulate it as the following matrix expression:
@@
    $\delta\boldsymbol{d}$ is a column vector where each element $\delta d_{j}$ represents the value of image residuals provided by the macro model for each image pixel, where $j=1,\ldots,N_{d}$, and $N_{d}$ denotes the number of image pixels. Similarly, the vector $\boldsymbol{n}$ represents the noise realisation at each image pixel, with values drawn from the multivariate normal distribution $\mathcal{N}(0,\,\mathrm{C}_{\mathrm{D}})$, where $\mathrm{C}_{\mathrm{D}}$ is the noise covariance matrix.
@@
    $\boldsymbol{B}$ is a square matrix of dimensions $N_{d},N_{d}]$ constructed to account for the blurring effect of the PSF.
@@
    $\boldsymbol{\beta_{j}^{d}}$ represents the ray-traced source plane position of the $j$-th native image grid (with position denoted by $\boldsymbol{\theta_{j}^{d}}$), according to the macro model. Let $S$ be the source brightness distribution predicted by the macro model. The $x$ and $y$ components of the source’s brightness gradient are given by $\frac{\partial S\left(\boldsymbol{\beta_{j}^{d}}\right)}{\partial\beta_{x}}$ and $\frac{\partial S\left(\boldsymbol{\beta_{j}^{d}}\right)}{\partial\beta_{y}}$.
@@
For convenience, we abbreviate $\boldsymbol{B}\boldsymbol{D}_{\mathrm{s}}\boldsymbol{C}_{\mathrm{f}}\boldsymbol{D}_{\psi}$ as $\mathbf{L_{\delta\psi}}$. The latter has dimensions of $N_{d}\times N_{p}$.
It is evident that $\delta\boldsymbol{d}$ and $\boldsymbol{\delta\psi}$ are connected via a linear mapping. Following the classical ‘semi-linear inversion’ framework of Warren and Dye (2003), the maximum a posteriori estimate for $\boldsymbol{\delta\psi}$, for a given $\mathbf{L}_{\delta\psi}$ and regularisation $\mathbf{R}_{\delta\psi}$, can be obtained via
$\boldsymbol{\delta\psi}_{\mathrm{MAP}}=\left(\mathbf{L}^{\mathrm{T}}_{\delta\psi}\mathbf{C}_{\mathrm{D}}^{-1}\mathbf{L}_{\delta\psi}+\mathbf{R}_{\delta\psi}\right)^{-1}\mathbf{L}^{\mathrm{T}}_{\delta\psi}\mathbf{C}_{\mathrm{D}}^{-1}\delta\boldsymbol{d}.$ (9)
Here, the regularisation matrix $\mathbf{R}_{\delta\psi}$ imposes a covariance prior on the pixelised potential correction $\boldsymbol{\delta\psi}$, thereby reducing the effective degrees of freedom and avoiding ill-posed solutions. The optimal regularisation $\mathbf{R}_{\delta\psi}$ can be determined by maximising the Bayesian evidence $\mathcal{E}$, whose formula was derived by Suyu et al. (2006) and reorganised by Dye et al. (2008), as follows:
$\log\mathcal{E}=$ $-\frac{N_{\mathrm{d}}}{2}\log(2\pi)-\frac{1}{2}\log(\det\mathbf{C}_{\mathrm{D}})$ (10)
$+\frac{1}{2}\log(\det\mathbf{R}_{\delta\psi})-\frac{1}{2}\delta\boldsymbol{\psi}^{\mathrm{T}}_{\mathrm{MAP}}\mathbf{R}_{\delta\psi}\delta\boldsymbol{\psi}_{\mathrm{MAP}}$
$-\frac{1}{2}(\mathbf{L}_{\delta\psi}\delta\boldsymbol{\psi}_{\mathrm{MAP}}-\delta\boldsymbol{d})^{\mathrm{T}}\mathbf{C}_{\mathrm{D}}^{-1}(\mathbf{L}_{\delta\psi}\delta\boldsymbol{\psi}_{\mathrm{MAP}}-\delta\boldsymbol{d})$
$-\frac{1}{2}\log[\det(\mathbf{L}^{\mathrm{T}}_{\delta\psi}\mathbf{C}_{\mathrm{D}}^{-1}\mathbf{L}_{\delta\psi}+\mathbf{R}_{\delta\psi})].$
A more thorough description of the full Bayesian framework can be found in Appendix A.